### v1

# Predicting and Understanding College Student Mental Health with Interpretable Machine Learning

Abstract.
Mental health issues among college students have reached critical levels, significantly impacting academic performance and overall wellbeing. Predicting and understanding mental health status among college students is challenging due to three main factors: the necessity for large-scale longitudinal datasets, the prevalence of black-box machine learning models lacking transparency, and the tendency of existing approaches to provide aggregated insights at the population level rather than individualized understanding.
To tackle these challenges, this paper presents I-HOPE, the first Interpretable Hierarchical mOdel for Personalized mEntal health prediction. I-HOPE is a two-stage hierarchical model, validated on the College Experience Study, the longest longitudinal mobile sensing dataset. This dataset spans five years and captures data from both pre-pandemic periods and the COVID-19 pandemic. I-HOPE connects raw behavioral features to mental health status through five defined behavioral categories as interaction labels. This approach achieves a prediction accuracy of 91%, significantly surpassing the 60-70% accuracy of baseline methods. In addition, our model distills complex patterns into interpretable and individualized insights, enabling the future development of tailored interventions and improving mental health support. The code is available at https://github.com/roycmeghna/I-HOPE.
Mental Health, College Students, Interpretable Machine Learning, Personalized Prediction
ccs: Computing methodologies Machine learning approachesccs: Human-centered computing Empirical studies in ubiquitous and mobile computingccs: Applied computing Health informaticsjournalyear: 2025copyright: ccconference: ACM/IEEE International Conference on Connected Health: Applications, Systems and Engineering Technologies; June 24–26, 2025; New York, NY, USAbooktitle: ACM/IEEE International Conference on Connected Health: Applications, Systems and Engineering Technologies (CHASE ’25), June 24–26, 2025, New York, NY, USAdoi: 10.1145/3721201.3721372isbn: 979-8-4007-1539-6/2025/06
1. Introduction
Mental health issues among college students have escalated to critical levels, significantly affecting academic performance, social interactions, and overall wellbeing (Hammoudi Halat et al., 2023; Chu et al., 2023; Barbayannis et al., 2022; Colarossi, 2022). The American College Health Association reports that 40% of students experience severe depression that disrupts daily functioning, while 60% encounter overwhelming anxiety during the 2020–2021 school year (Abrams, 2022). Furthermore, approximately 76% of college students report moderate to severe psychological distress, with anxiety and depression being the most prevalent diagnoses (BestColleges, 2024). Despite increasing awareness, timely access to support remains limited for many students due to stigma, resource constraints, and challenges in detecting those at risk (World Health Organization, 2024). This situation underscores the pressing need for effective and scalable solutions to improve the understanding and prediction of mental health outcomes.
The complexity of predicting and understanding mental health status among college students arises from three primary factors. First, a comprehensive analysis requires a large-scale, longitudinal dataset that collects data through passive sensing over an extended period rather than relying on short-term data collection conducted in a lab setting. Second, although machine learning has been used to address mental health issues, many existing models utilize black-box algorithms that lack transparency and interpretability (Xu and Shuttleworth, 2024; Sajno et al., 2023; Tutun et al., 2023). Third, most machine learning approaches yield aggregated insights at the population level that fail to provide individualized understanding, which is essential for personalized interventions and mental health support (Batstra and Timimi, 2024; Conley, 2024).
In this paper, we address these challenges through a paradigm shift towards methodologies that not only leverage an extensive dataset but also prioritize individual variability in mental health prediction. In particular, we leverage the College Experience Study (CES) dataset (Nepal et al., 2024), the longest longitudinal mobile sensing dataset for college student behaviors, released by Dartmouth College in October 2024. This dataset is especially valuable for our study because it covers a five-year period that includes pre-pandemic years, the COVID-19 pandemic, and the gradual return to normalcy as the pandemic receded. By analyzing behavioral patterns and mental health metrics over different time periods, we can investigate insights into the pandemic’s impact on mental health and the role of behavior in shaping it.
We present I-HOPE, the first Interpretable Hierarchical mOdel for Personalized mEntal health prediction. I-HOPE is a two-stage hierarchical model designed to accurately predict mental health status while offering deep insights into the features contributing to various mental health conditions. Specifically, we define five interaction labels: Leisure, Me Time, Phone Time, Sleep, and Social Time, which categorize different daily behaviors. The key insight of I-HOPE is to connect raw behavioral features to mental health status through an intermediate layer consisting of these five interaction labels. These labels act as compact representations of complex behaviors, simplifying the input space while preserving data richness. Furthermore, I-HOPE facilitates the identification of an individual’s social type and emphasizes the interactions that most significantly impact their mental health.
We compare I-HOPE with baseline methods that lack personalized predictions and two-stage feature mappings. I-HOPE achieves an overall prediction accuracy of 91%, significantly surpassing the 60-70% accuracy of baseline methods. I-HOPE simplifies complex patterns into interpretable and personalized insights by mapping behavioral features into interaction labels. For instance, a good 7-hour sleep is linked to better mental health outcomes, while poor sleep correlates with increased anxiety and depression. Walking serves as a stress reliever associated with leisure and relaxation. Phone usage reveals contrasting patterns; interactions at home often indicate positive connections, whereas excessive use in social settings suggests stress or discomfort. The balance between social and personal time is crucial, with shared spaces that promote engagement and personal spaces that allow emotional recharge. Importantly, these behaviors affect individuals differently; some find “Me Time” in working out, while others prefer studying or spending time in their own dorms.
We summarize the contributions as follows:
  • Interpretable machine learning framework development: We present I-HOPE, the first hierarchical model that maps raw behavioral data into five interpretable interaction labels: Leisure, Me Time, Phone Time, Sleep, and Social Time, enhancing the transparency and accuracy of predictions.
  • Personalized mental health predictions: I-HOPE adapts to individual behaviors, enhancing the accuracy and relevance of mental health assessments.
  • Key behavioral predictors identification: I-HOPE identifies specific behaviors, such as sleep patterns and physical activity levels, significantly influencing mental health outcomes, providing insight into targeted interventions.
  • Scalable analytical approach: I-HOPE scales high-dimensional behavioral data analysis, linking complex datasets to real-world mental health applications.
2. College Experience Study (CES) Dataset
To thoroughly investigate college students’ mental health status, a large-scale dataset that captures daily behaviors is required. Therefore, we leverage the CES dataset (Nepal et al., 2024), the longest longitudinal mobile sensing dataset for college student behaviors, released by Dartmouth College in October 2024. This dataset includes passive mobile sensing data— mobility, physical activity, sleep patterns, and phone usage—along with Ecological Momentary Assessment (EMA) surveys from 217 Dartmouth students collected between 2017 and 2022. It comprises over 210,000 data points collected across two cohorts throughout their college years on an hourly basis. The EMA surveys are delivered randomly once a week via the StudentLife mobile application (Wang et al., 2014).
This dataset is especially valuable for our study because it covers a five-year period that includes pre-pandemic years, the COVID-19 pandemic, and the gradual return to normalcy as the pandemic receded. By analyzing behavioral patterns and mental health metrics over different time periods, we can assess and predict how mobile sensing data—collected before, during, and after the COVID-19 pandemic—affects students’ mental health. This analysis provides valuable insights into the pandemic’s impact on mental health and the role of behavior in shaping it.
Table 1. PHQ-4 scores and their categories.
PHQ-4 score Category
0-3 Normal
4-6 Mild
7-9 Moderate
10-12 Severe
PHQ-4 score.
The EMA survey includes a key mental health metric, Patient Health Questionnaire-4 (PHQ-4) score (Kroenke et al., 2009), which serves as the focal point for our analysis and the mental health outcome we aim to predict in this paper. The PHQ-4 score is a widely recognized screening tool for assessing depressive and anxiety symptoms in clinical settings and epidemiological studies (Löwe et al., 2010). It ranges from 0 to 12, with lower scores indicating better mental health. To enhance interpretability and align with standardization practices (Wicke et al., 2022), we categorize PHQ-4 scores into four levels: normal, mild, moderate, and severe. Each category corresponds to specific PHQ-4 score ranges as shown in Table 1, which allows clearer insight into mental health status. The distribution of data points across the PHQ-4 categories is shown in Figure 1.
Figure 1. Imbalanced data distribution across PHQ-4 categories. PHQ-4 category Normal has the highest count, whereas PHQ-4 category Severe has the lowest.
Features.
The dataset includes 172 features, including smartphone activity, duration spent performing various activities, duration spent at various locations, sleep information, etc. However, not all features directly relate to mental health studies, such as the amplitude of detected audio. Additionally, some features, such as distance traveled and number of locations visited, are ambiguous. After careful screening, we select 35 features for our study. These features, which include contextual information, phone usage, and sleep duration, are chosen based on the following criteria.
  1. (1)
    They provide a meaningful representation of mental health status by capturing behavioral patterns that research has consistently linked to mental health.
  2. (2)
    They demonstrated statistical significance with respect to the PHQ-4 categories, with p-values less than 0.05 and, in many cases, less than 0.01.
3. Motivation
Table 2. Pearson correlations between PHQ-4 score with smartphone usage features in five environments.
Category Metric Correlation
Overall Duration of Unlock 0.0321
# Unlocks -0.0107
Food Duration of Unlock -0.0125
# Unlocks -0.0197
Study Duration of Unlock 0.0014
# Unlocks -0.0151
Social Duration of Unlock -0.01
# Unlocks -0.0157
Dormitory Duration of Unlock -0.00215
# Unlocks -0.01505
Home Duration of Unlock 0.0188
# Unlocks 0.0033
Table 3. Interaction labels and their corresponding relevant features.
Interaction label Relevant features corresponding to each interaction label
Leisure Duration of physical activities (biking, walking, running) Duration of conversations (in-person and phone) Duration in locations like others’ dorm, workout Duration of phone usage in various locations
Me Time Duration of activities (biking, walking, running, being still, studying) Various measures at home (conversation detected, duration) Conversations detected at own dorm or home Duration of phone usage in various locations
Phone Time Duration of phone conversations in various locations Ratio of number of calls to duration of calls Ratio of number of phone unlocks to duration of unlocks, in various locations
Sleep Duration of stillness Audio detection at home or own dorm Phone usage at night Sleep duration
Social Time Duration of social activities (biking, walking, running, workout, study, eating food) Ratio of number of calls to duration of calls at locations like study space, home, etc. Amount of time spent in others’ dorm
In this section, we explore the challenges of predicting mental health status by analyzing the dataset and identify key insights for effective solutions.
3.1. The Role of Smartphone Usage on Mental Health
In mental health research, smartphone usage has increasingly been used to facilitate real-time data collection, improve access to mental health resources, and enable personalized interventions. Prior work has shown that screen time and social media use negatively impact mental health and digital wellbeing (Zhao et al., 2024; Elamin et al., 2024). Mobile sensing and interventions have also shown promise for personalizing mental health support (Yu et al., 2024). Therefore, we start by focusing on two smartphone usage metrics: the number of times phones are unlocked (#Unlocks) and daily phone usage duration (Duration of Unlock).
To explore the relationship between these features and mental health, we conduct a correlation analysis using the CES dataset. The results, summarized in Table 2, indicate three main findings: (i) longer phone usage correlates with poorer mental health; (ii) higher unlock frequencies are associated with better mental health; (iii) increased phone use in social and dining settings relates to worse mental health, possibly due to social anxiety, while greater use at home suggests positive activities like socialization, leisure, or learning. These findings highlight the importance of context in understanding the impact of smartphone usage on mental health.
3.2. Smartphone Usage Features Alone Are Insufficient for Predicting Mental Health
To assess whether smartphone usage features can effectively predict mental health outcomes, we train a machine learning model on the CES dataset to predict the PHQ-4 score. Given that the PHQ-4 score has four categories, as illustrated in Table 1, this constitutes a multi-class classification problem. Therefore, we use two smartphone features mentioned above—the number of unlocks and daily phone usage duration—as input features to train a multilayer perceptron (MLP) model, a neural network widely used for classification and prediction (Popescu et al., 2009). We preprocess the data by normalizing the features and ensuring balanced class distributions through oversampling for the PHQ-4 categories. We randomly split the dataset into 80% for training and 20% for testing. Our results reveal that the prediction accuracy is only 28%, which is low. Even after oversampling to achieve balanced class distributions for the PHQ-4 categories, the accuracy rises only to 30%. These results indicate that relying solely on these smartphone usage features is insufficient to predict mental health outcomes accurately.
3.3. Improving Predictions with Additional Features
To improve prediction accuracy, we need to incorporate additional features from the CES dataset. Specifically, we add 33 more features from the CES dataset related to location information, sleep duration, and activity times (see Column 2 of Table 3). With increased input dimensions, we use the same MLP model with 35 input nodes, three fully connected hidden layers, and four output nodes corresponding to the PHQ-4 categories. The training and test sets are the same as above. The model is trained for 50 epochs until the learning curve stabilizes. This approach improves the prediction accuracy to 60%, which is still not high. We attribute the low accuracy to two reasons: (i) Treating the dataset as a whole may have obscured individual differences; (ii) A high correlation among features likely reduce the model’s ability to make accurate predictions. We discuss these below:
Addressing Individual Differences
The CES dataset includes data from 217 individuals, each contributing about 160 data points. Research indicates that individual differences in behavior and preferences can significantly affect predictive modeling outcomes (He et al., 2023; Lengerich et al., 2019). To capture these variations, we develop personalized models by grouping data based on each student’s unique user ID (UID). Each model uses the same 35 features but is trained on the corresponding individual’s data, maintaining the original neural network architecture and train-test split. This personalized approach improves the average prediction accuracy to 70%. However, high feature correlation remains a challenge, introducing redundancy and limiting further accuracy improvements, highlighting the need for additional refinement to address these feature-related issues effectively.
Reducing Feature Correlation
Figure 2. Heatmap for correlations among features.
Highly correlated features introduce redundancy, leading to overfitting and reduced test accuracy while increasing computational complexity without adding valuable information. As shown in Figure 2, many dataset features are highly correlated. To address this, we analyze feature importance using random forest, an efficient and interpretable ensemble learning technique (Breiman, 2001). This method builds multiple decision trees and combines their outputs to evaluate feature importance. Figure 3 shows the importance ranking of all 35 features, highlighting the top ones. We select the top 50% of these features to retrain personalized neural network models, adjusting the input dimensions accordingly. However, this results in a drop in prediction accuracy to 65%, indicating that simply excluding certain features may remove critical information necessary for effective predictions.
Figure 3. Feature importances using random forests.
Key takeaway.
These analyses suggest that the features influencing prediction accuracy vary among individuals. For instance, a higher biking time may correlate with better mental health for one person, while walking might be more relevant for another, and a combination of running and walking could be key for someone else. A global feature importance approach risks eliminating crucial features for individuals, potentially reducing accuracy. Therefore, it is vital to incorporate personalization while minimizing feature correlation, ensuring the model captures individual-specific patterns without being hindered by redundant or irrelevant features, ultimately enhancing overall prediction accuracy.
Figure 4. Workflow of the I-HOPE design.
4. I-HOPE
So far, our analysis has indicated that globally selecting features for model training reduces correlation but does not enhance prediction accuracy. This underscores the need for a personalized approach that minimizes feature correlation while capturing individual-specific patterns, improving prediction accuracy and interpretability.
Therefore, we design and implement I-HOPE, the first Interpretable Hierarchical mOdel for Personalized mEntal health prediction. I-HOPE is a two-stage hierarchical model that connects behavioral features to mental health status through an intermediate layer of five interaction labels: Leisure, Me Time, Phone Time, Sleep, and Social Time, which categorize different daily behaviors. This structure offers two benefits:
  1. (1)
    Dimension Reduction. The interaction labels serve as compact representations of complex behaviors, simplifying the input space while retaining the richness of the data.
  2. (2)
    Enhanced Interpretability. I-HOPE enables the identification of an individual’s social type and highlights the interactions that influence their mental health the most.
Figure 4 outlines the workflow of I-HOPE design. It starts with n raw input features ($n=35$ from the CES dataset), which are mapped to k interaction labels ($k=5$), representing aggregated behavioral categories: Leisure, Me Time, Phone Time, Sleep, and Social Time. To quantify the behavioral insights, we define a score for each label and use the scores as input for Stage 2. I-HOPE uses a thresholding mechanism that incorporates personalized feature importance to calculate interaction label scores, effectively reducing feature correlation by condensing the 35 raw features into five labels. These scores serve as input to a neural network that predicts one of four mental health categories based on the PHQ-4 scale: Normal, Mild, Moderate, or Severe. I-HOPE groups individual preferences under interaction labels, improving predictive consistency. It also improves interpretability by identifying which interaction label score influences the predictions and determining the significance of specific raw features within each label. For example, I-HOPE can assess whether “walking” or “spending time with friends” in the Leisure category has a greater impact on mental health of an individual. Next, we will describe each stage of I-HOPE and explain the threshold mechanism that links them.
4.1. Stage 1: Input Features $\rightarrow$ Interaction Labels
In Stage 1, I-HOPE maps raw input features to five interaction labels: Leisure, Me Time, Phone Time, Sleep, and Social Time. These labels group features with statistically significant associations to mental health outcomes (p¡0.05 or p¡0.01) into each label. This approach combines domain knowledge and data-driven insights to capture relevant behavioral dimensions, reduce redundancy, and enhance interpretability while maintaining granularity for personalized analysis. The mapping process is described below.
  1. (1)
    Feature selection and initial mapping: We manually select a subset of features from the dataset ($n=35$) to match each interaction label based on their semantic and behavioral relevance. For instance, features like unlock duration and phone conversations belong to Phone Time, while duration of social activities and time spent with friends belong to Social Time. The initial mapping is detailed in Table 3. Notably, some features are assigned to multiple labels, reflecting their multidimensional impact on various behavioral categories.
  2. (2)
    Feature clustering: To group features with similar behavioral patterns and ensure consistency within interaction labels, I-HOPE uses K-means clustering to identify inherent groupings that capture meaningful behavioral dimensions without predefined ground truth.
  3. (3)
    Feature importance analysis: I-HOPE uses a random forest model to assess the contribution of individual input features to each interaction label, as it can handle high-dimensional data and capture nonlinear relationships (Breiman, 2001). Random forest measures feature importance by evaluating how much each feature reduces uncertainty at decision splits across all trees. Features that significantly improve data separation receive higher importance scores, which are then used to refine interaction label scoring. This analysis is conducted per-user basis, grouping data by UIDs.
Algorithm 1 Calculation of LeisureScore
1:  Input: Threshold of features mapped to Leisure across UIDs , Feature Importances (FeatImps)
2:  Initialize: LeisureScore = 0
3:  if Walking Duration $<$ Threshold (where Threshold = Mean of distribution) then
4:     LeisureScore += FeatImp of “Walking Duration”
5:  end if
6:  if Phone usage at home $<$Threshold (where Threshold = Mean of distribution) then
7:     LeisureScore += FeatImp of “Phone usage at home”
8:  end if
9:  if Running Duration $<$ Threshold (where Threshold = Mean of distribution) then
10:     LeisureScore += FeatImp of “Running Duration”
11:  end if
12:  if Duration of Conversation $>$ Threshold (where Threshold = Mean of distribution) then
13:     LeisureScore += FeatImp of “Duration of Conversation”
14:  end if
15:  if Phone usage at others’ dorm $<$ Threshold (where Threshold = Mean of distribution) then
16:     LeisureScore += FeatImp of “Phone usage at others’ dorm”
17:  end if
18:  if Voice Detection in others’ dorm $>$ Threshold (where Threshold = Mean of distribution) then
19:     LeisureScore += FeatImp of “Voice Detection in others’ dorm”
20:  end if
21:  if Duration of conversation in others’ dorms $>$ Threshold (where Threshold = Mean of distribution) then
22:     LeisureScore += FeatImp of Duration of conversation in others’ dorms
23:  end if
Algorithm 2 Calculation of SleepScore
1:  Input: Threshold of features mapped to label Sleep across UIDs, Feature Importances (FeatImps)
2:  Initialize: SleepScore = 0
3:  if Sleep duration $>$ Threshold (where Threshold = Mean of distribution) then
4:     SleepScore += FeatImp of “Sleep duration”
5:  end if
6:  if Duration at own dorm $>$ Threshold (where Threshold = Mean of distribution) then
7:     SleepScore += FeatImp of “Duration at own dorm”
8:  end if
9:  if Duration of stillness $>$ Threshold (where Threshold = Mean of distribution) then
10:     SleepScore += FeatImp of “Duration of stillness”
11:  end if
12:  if Conversation detected at home $<$ Threshold (where Threshold = Mean of distribution) then
13:     SleepScore += FeatImp of “Conversation detected at home”
14:  end if
4.2. Threshold based Scoring
This section outlines the scoring mechanism for interaction labels. Each label captures relevant behavioral information, allowing us to assess the importance of features for individuals. Since these labels are qualitative, they must be converted into quantitative features for predictive analysis. To achieve this, we define a score, termed [InteractionLabel]Score, for each label based on thresholds derived from the distribution of associated characteristics. The score is updated by checking if a feature’s value exceeds or falls below the threshold and adjusting it according to its importance.
To ensure that the scores are tailored to individual users, I-HOPE updates the interaction label scores by assigning greater weight to features with higher importance scores. This personalized approach enables a more accurate representation of each label and enhances the relevance of the features for predictive modeling. Each label’s score is the sum of feature contributions that meet the threshold, where contributions are weighted based on feature importance. These scores are then used as input features for predictive modeling.
To demonstrate this methodology, we use the labels Leisure and Sleep as examples. These labels are selected for clarity and simplicity, but the same process applies uniformly to all other labels.
  • The LeisureScore is derived using the following features (Algorithm 1): duration of walking, ratio of phone unlocks to duration of phone use, duration of running, duration of conversation, ratio of phone unlocks to duration of phone use in others’ dorms, and duration of conversation in others’ dorms.
  • The SleepScore is derived using the following features (Algorithm 2): sleep duration, duration of the phone being idle, time spent in own dorm, and duration of conversation detected.
4.3. Stage 2: Interaction Labels $\rightarrow$ PHQ-4 Category
In Stage 2, I-HOPE uses the five interaction label scores (LeisureScore, MeScore, PhoneScore, SleepScore, SocialScore) as inputs to predict PHQ-4 categories. This stage simplifies the predictive process by using these labels as compact and interpretable representations of the raw data. The process includes the following steps:
  1. (1)
    Neural network prediction: I-HOPE builds a neural network model to predict PHQ-4 categories using the five interaction label scores as inputs. The neural network includes an input layer with five nodes, three hidden layers for pattern extraction, and an output layer with four nodes for the PHQ-4 categories (normal, mild, moderate, and severe). I-HOPE uses the ReLU activation function for hidden layers and the Softmax activation function for the output layer for multi-class classification. The neural network model is trained for 50 epochs using the Adam optimizer (Kingma and Ba, 2014) with a learning rate of 0.001 and categorical cross-entropy as the loss function.
  2. (2)
    Weighted feature representation: The input scores of the neural network are weighted by the importance of the feature of the random forest analysis, ensuring that the model captures individual variations and enhances the accuracy of prediction.
  3. (3)
    Performance evaluation: I-HOPE calculates the mean test accuracy across all UIDs to measure performance while accounting for individual differences. Using interaction labels reduces input complexity, highlighting the effectiveness of I-HOPE’s hierarchical design.
5. Evaluation
In this section, we evaluate the I-HOPE to address the following research questions:
  1. RQ1.
    Prediction accuracy: How accurately does I-HOPE predict the mental health status of college students?
  2. RQ2.
    Interpretability from hierarchical mapping: What new insights into mental health can we gain from hierarchical feature mapping of I-HOPE?
  3. RQ3.
    Scoring for interaction labels: How much do features contribute to the scoring of interaction labels?
To answer these questions, we used Python tools for data preprocessing (cleaning, transformation, and feature extraction) with Pandas and NumPy. I-HOPE was implemented, trained and tested using TensorFlow, while evaluation metrics such as accuracy and loss are computed with Scikit-learn. We randomly split the dataset into 80% for training and 20% for testing.
The rest of this section is divided into three parts, each addressing a research question with detailed results and insights.
5.1. Prediction Accuracy
Table 4. Comparison of different methods for predicting PHQ-4 categories. Evaluation metrics include precision (Pre.), recall (Rec.), F1-score (F1), and overall accuracy for each PHQ-4 category.
PHQ-4 Category Baseline 1 Baseline 2 Baseline 3 I-HOPE
Prec. Rec. F1 Prec. Rec. F1 Prec. Rec. F1 Prec. Rec. F1
Normal 0.6453 0.627 0.6360 0.7149 0.677 0.6954 0.6781 0.654 0.6659 0.9513 0.9285 0.9398
Mild 0.6135 0.617 0.6152 0.6432 0.667 0.6549 0.6297 0.667 0.6478 0.9489 0.9285 0.9386
Moderate 0.5908 0.600 0.5953 0.6371 0.660 0.6483 0.6083 0.660 0.6332 0.8791 0.8690 0.8740
Severe 0.5831 0.559 0.5708 0.6265 0.614 0.6202 0.5982 0.614 0.6060 0.8645 0.9138 0.8884
Overall Accuracy 0.60 0.70 0.65 0.91
Figure 5. Normalized confusion matrix showing high classification accuracy across all categories from I-HOPE.
We compare I-HOPE against several baseline methods that differ in the training approach (aggregated vs. personalized), feature extraction (all raw features vs. selected features) and model structure (single stage vs. hierarchical). The methods are summarized as follows, all using the previously described neural network architecture with adjusted input dimensions:
  • Baseline 1: A single model trained with all raw features.
  • Baseline 2: Personalized models for each individual using all raw features.
  • Baseline 3: Personalized models for each individual using the top 50% of features based on feature importance from random forest models.
  • I-HOPE: The Interpretable Hierarchical mOdel for Personalized mEntal health prediction presented in this paper.
The evaluation results over different metrics are summarized in Table 4. Baseline 1 achieves only 60% accuracy due to feature redundancy and lack of personalization. Baseline 2 improves to 70% but still faces feature correlation and redundancy issues. Baseline 3 reduces feature correlation using feature importance but drops to 65% accuracy by excluding significant features. In contrast, I-HOPE reaches 91% overall prediction accuracy by using interaction labels to capture distinct user behaviors while balancing interpretability and predictive power. We can see that the baseline methods struggle with feature redundancy and adapting to user-specific patterns, while I-HOPE effectively integrates personalized interaction labels and feature importance, showing better precision, recall and hence F1 score too. We further use the confusion matrix to provide a detailed breakdown of I-HOPE’s predictions against the true labels. As shown in Figure 5, I-HOPE accurately classifies a majority (more than 90% in three cases) of data points across all four categories. Overall, the evaluation using multiple metrics highlights the strong predictive performance of I-HOPE.
5.2. Interpretability
To gain insights into mental health and improve prediction accuracy, it is crucial to extract meaningful and interpretable features from raw behavioral data. I-HOPE simplifies and organizes raw data into actionable constructs, enabling the capture of individual behavioral patterns and their connection to mental health outcomes.
Figure 6. The individual-level heatmaps that highlight feature importance for each interaction label across 217 UIDs. The y-axis represents label-specific features, which vary across different interaction labels.
5.2.1. Evaluation of Stage 1: Input Features $\rightarrow$ Interaction Labels
The first stage of I-HOPE maps the raw input features into five interaction labels: Leisure, Me Time, Phone Time, Sleep, and Social Time. This simplifies the dataset while preserving key behavioral patterns relevant to mental health. To enhance interpretability, we conduct feature importance analysis for all selected features within each label. This analysis identifies which features most influence the model’s predictions, revealing meaningful behavioral patterns related to mental health. We use a random forest model to compute feature importance scores, providing an interpretable ranking of features based on their contributions. This analysis is performed individually for personalized insights into user behavior.
Figure 6 shows personalized feature importance across five labels using heatmaps for 217 students. The analysis reveals distinct patterns of feature relevance for each label, highlighting both shared and unique predictors. Below, we summarize the key findings and insights for each label.
  • Label: Leisure - Feature F4 (Duration of walking) emerged as the most important predictor for 90.36% of UIDs, emphasizing walking as a key leisure activity. Other features, such as F13 (Phone usage at home), F3 (Duration of running), F5 (Duration of conversations), F8 (Voice detection at others’ dorms), and F14 (Phone usage at others’ dorms), are moderately important for 20–25% of UIDs, indicating that leisure also encompasses social interactions and phone use. In contrast, features like F1 (Biking duration), F2 (Footsteps), F6 (Duration of calls), F9 (Conversations at others’ dorms), F10 (Audio detections at social places), F11 (Duration at “leisure”), and F12 (Workout duration) consistently show lower relevance. F11 has low importance despite its direct link to the label, likely due to unclear dataset definitions, underscoring the need for improved data clarity. These findings illustrate how behaviors like walking and social interactions influence leisure and provide insights into broader themes related to mental health and lifestyle habits.
  • Label: Me Time - Features such as F16 (Phone usage at own dorm), F12 (Time spent at own dorm), F7 (Voice detections at own dorm), F13 (Study duration), and F17 (Phone usage at study locations) are important for 89.5% UIDs, with F12 and F16 being the most significant. These findings suggest that activities in dorm spaces, such as studying and socializing, are key indicators of “Me Time” and contribute to a supportive environment. Conversely, features like F1 (Biking duration), F2 (Footsteps), F3 (Running duration), F5 (Walking duration), and F14 (Workout duration) show low importance, indicating that individual physical activities are less relevant to this label. Other features display varying levels of significance across individuals, reflecting the diverse experiences of “Me Time”. These results underscore the role of personal environments like dorms in fostering meaningful routines and behaviors, offering insights into how individuals balance solitude, productivity, and emotional wellbeing.
  • Label: Sleep - Features F6 (Sleep duration) and F7 (Duration of being idle) are key predictors, highlighting the importance of inactivity and sleep in identifying sleep patterns. Features like F1 (Conversation detection at dorm) and F4 (Location at dorm) show moderate importance, suggesting that reduced social interactions and staying in the dorm contribute to recognizing sleep behaviors, likely linked to quiet nighttime routines. In contrast, features such as F2 (Voice detections at home), F3 (Conversations at home), F5 (Phone usage at home), and F8 (Health fitness) are minimally relevant. These results emphasize the role of inactivity and dorm contexts in understanding sleep patterns while activities like fitness or phone use have a limited impact on sleep.
  • Label: Social Time - Feature F15 (Duration at study location) is the key predictor, indicating a strong link between study locations and social activities. Other features include F2 (Walking duration), F7 (Voice detections at study areas), F13 (Time spent in others’ dorms), and F16 (Phone usage in others’ dorms), which highlight the importance of mobility, conversations, and phone use. In contrast, features like F1 (Footsteps), F9 (Conversations in others’ dorms), F10 (Conversations at social places), and F12 (Duration of leisure) show limited importance, suggesting they are less directly related to social time or relevant only for a few individuals. These findings emphasize the role of specific environments and interaction patterns in shaping social time while indicating that factors like leisure duration or general mobility have a minor impact.
  • Label: Phone Time - Feature F10 (Phone usage at own dorm) is the primary predictor, indicating that students primarily use their phones in personal spaces where they feel comfortable engaging in calls and messaging. Other features include F5 (Audio detected at own dorm), F7 (Audio detected at home), F8 (Phone usage at home), F12 (Phone usage at study locations), and F13 (Phone usage throughout the day), which highlight variations in phone use across personal, home, and study contexts. Conversely, features like F1 (Audio conversations detected throughout the day), F2 (Number of SMS), F3 (Voice detection at home), F4 (Voice detection in social settings), F6 (Conversations detected at home), F9 (Phone usage in others’ dorms), and F10 (Phone usage in workout areas) show minimal relevance, likely due to ambiguous definitions or differences in data collection between iOS and Android devices. These findings emphasize the role of personal and study spaces in shaping phone usage while highlighting the need for clearer feature definitions and standardized data collection to enhance the reliability of less relevant features.
In summary, the findings reveal personalized variability among students in feature importance, with key behavioral patterns linked to mental health and activity contexts:
  • Behavioral patterns: Walking, phone usage, and time in personal spaces are key to labels like Leisure, Phone Time, and Me Time. Walking supports relaxation, while dorm-based phone use reflects connection and self-reflection, linking physical spaces to emotional wellbeing.
  • Activity contexts: Study locations and others’ dorms strongly predict Social Time, emphasizing collaborative settings. In contrast, private spaces dominate Me Time and Phone Time, illustrating the balance between shared and personal spaces for social engagement and self-care.
  • Sleep and rest: Sleep duration and idle time are central to Sleep, with dorm-related contexts suggesting restful routines tied to spatial habits and reduced interactions, reinforcing sleep’s role in mental health.
  • Personalized variability as a key factor: The varying importance of features like walking or conversation detection reflects the individual nature of behavior. Tailored analyses can enhance mental health predictions, making them more accurate and context-aware.
These results emphasize the importance of selecting features that capture distinct behavioral dimensions for interaction labels. They also demonstrate that personalized analysis is key for accurate mental health prediction, as feature importance varies across users. This forms the basis for our label scoring and predictions, explained in the following subsections.
Figure 7. The individual-level heatmap that shows the importance of interaction label scores in predicting PHQ-4 categories. Each row represents an interaction label across 217 individuals.
5.2.2. Evaluation of Stage 2: Interaction Labels $\rightarrow$ PHQ-4 Categories
Figure 7 shows the importance of interaction labels (Leisure, Me Time, Phone Time, Sleep, Social Time) in predicting PHQ-4 mental health categories across 217 users. Each row corresponds to a label, and each column represents an individual user, with colors ranging from dark purple (low importance) to bright yellow (high importance). The figure highlights how the relevance of these labels varies significantly across individuals, revealing the personalized nature of mental health predictors.
Among the labels, Sleep is the most consistently important, with high relevance for 95% of students, reflecting the strong link between healthy sleep patterns and stable mental health. Phone Time also shows significant importance across many users, likely indicating stress or coping behaviors related to screen use. In contrast, labels like Social Time and Leisure exhibit more variability; Social Time is crucial for some but less so for others, reflecting individual differences in social interactions’ impact on mental health. Similarly, Leisure is an important predictor for a subset of users, emphasizing the role of relaxation activities. However, Me Time serves as a predictor for an even smaller group, highlighting the challenges of finding solitude in a school environment with shared spaces.
In all, Figure 7 underscores a critical insight – mental health predictors are not one-size-fits-all. Each label contributes uniquely, reflecting the diverse and personalized ways behavioral patterns influence mental health outcomes. While Sleep and Phone Time act as general predictors across most users, Social Time, Me Time, and Leisure provide insights tailored to specific individuals. By combining these labels, we can build a comprehensive understanding of behavioral patterns and their impact on mental health, supporting the need for personalized modeling approaches to predict PHQ-4 categories effectively.
Figure 8. Mean feature importance across all UIDs for each interaction label.
5.3. Scoring for Interaction Labels
To analyze how each feature contributes to scoring for interaction labels, we visualize the importance of relevant features across all 217 students. Figure 8 displays the top features (marked with a star) for each label:
  • Leisure: F3 (running duration), F4 (walking duration) , F5 (conversation duration), F8 (voice detection at others’ dorms), F13 (phone usage at home), and F14 (phone usage at others’ dorms).
  • Me Time: F7 (voice detections at own dorm), F12 (time spent in own dorm), F13 (study duration), F16 (phone usage at own dorm), and F17 (phone usage at study locations).
  • Sleep: F1 (conversations detected at dorm), F4 (duration at own dorm), F6 (sleep duration), and F7 (duration of being idle).
  • Social Time: F2 (walking duration), F7 (voice detected at study location), F13 (time in others’ dorms), F15 (duration at study location), and F16 (phone usage at others’ dorm).
  • Phone Time: F5 (audio detected at own dorm), F8 (phone usage at home), F10 (phone usage at own dorm), F12 (phone usage at study locations), and F13 (phone usage throughout the day). Phone usage corresponds to the ratio of the number of phone unlocks to usage duration.
The scores for each interaction label are calculated based on the relative importance of various features, measured against a threshold value derived from the mean distribution of each feature across all UIDs. The scoring process increases the label’s score based on the contribution of features that meet the threshold condition, as outlined in Section 4.2.
Figure 9. Thresholds for LeisureScore calculation based on six features. The vertical dashed red line is the threshold.
Figure 9 and Figure 10 depict the thresholds (represented as the mean values of the distributions, shown by the dashed red lines) for each feature within their respective interaction labels. In Figure 9, the dotted line highlights the threshold values for the most important features (F4, F13, F3, F5, F14, and F8). F4 reflects walking duration: if a student’s walking duration is less than the threshold (3 hours), it indicates leisure activity. Consequently, the LeisureScore is incremented proportionally to the feature importance of walking for the respective student. Similarly, F8 considers voice detection in others’ dormitories. If the duration exceeds the threshold (15 minutes), it suggests that the student is having leisure time with friends, causing an increase in LeisureScore by the value of importance of the corresponding feature. The same logic is applied to the other features, as shown in the figure and detailed further in Algorithm 1.
In Figure 10, the thresholds for the most important features (F6, F4, F7, and F1) are delineated by the dotted line. F6 indicates total sleep duration: if a student’s sleep exceeds 7 hours, representing the standard benchmark for healthy sleep, the SleepScore is incremented based on the student’s feature importance for sleep. F4, which measures the time spent in the room, shows that durations greater than 12.9 hours typically signify overnight dorm occupancy, often indicative of sleep rather than other activities. F7 corroborates healthy sleep with thresholds above 7.5 hours, while F1 reflects quietness, with values below 6 minutes suggesting an environment conducive to rest. Each scenario contributes to the SleepScore as outlined in Algorithm 2.
Figure 10. Thresholds for SleepScore calculation based on four features. The vertical dashed red line is the threshold.
6. Related Work
Machine learning increasingly tackles various challenges in mental health research (Shatte et al., 2019; D’Alfonso, 2020; Kim et al., 2021; Sajno et al., 2023; Tutun et al., 2023; Adler et al., 2022). Using data from mobile devices that monitor smartphone use, social media activity, and medical records, researchers have developed machine learning techniques to identify and predict conditions such as depression, anxiety, and stress (Saeb et al., 2015; Guntuku et al., 2017; Zhao et al., 2024). Furthermore, natural language processing has opened new pathways for analyzing text data, including social media posts and therapy transcripts, offering valuable insights into emotional and psychological states (Eichstaedt et al., 2018; Althoff et al., 2016; Cook et al., 2016; Stewart and Velupillai, 2021; Lamichhane, 2023).
However, despite these advances, many machine learning techniques struggle to capture and understand mental health characteristics at the individual level. Researchers often focus on applying techniques without addressing the issues of variability, personalization, and interpretability (Mohr et al., 2017; Torous et al., 2016; Batstra and Timimi, 2024; Moggia et al., 2024). This limitation restricts the practical application of these methods, particularly in treatment recommendations and interventions.
The release of the College Experience Study dataset offers a unique opportunity to analyze the longest longitudinal passive mobile sensing dataset on the mental health of college students at the individual level (Nepal et al., 2024). Our prior work was the first attempt to analyze this dataset (Xuan et al., 2025). This paper goes a step further by addressing a key research gap with I-HOPE, the first interpretable hierarchical model designed to accurately predict and offer new insights into the mental health of college students.
7. Conclusion
This paper addresses the challenges of predicting and understanding mental health among college students. Using a cutting-edge longitudinal dataset focused on college student behavior, we present I-HOPE, the first hierarchical model designed to predict and interpret mental health status at the individual level. Our evaluation shows that I-HOPE significantly outperforms baseline methods, offering valuable insights into how various behavioral features influence mental health outcomes. This research lays the foundation for future efforts aimed at predicting and understanding the mental health of college students. We hope that our findings inspire further exploration and innovation in this critical area of study, ultimately contributing to enhanced support systems for student wellness.
Acknowledgements
We thank anonymous reviewers for their helpful feedback to improve the final version of the paper. The work of Meghna Roy Chowdhury and Shreyas Sen is supported by the National Science Foundation under Grant # 1944602.